OpenStudy (anonymous):

Jake Bass bought a home for $170,000 putting down $40,000. The rate of interest is 14 percent for 25 years. The total yearly mortgage payment is

OpenStudy (anonymous):

8 percent of 450

OpenStudy (anonymous):

\[\frac{130,000\times \frac{.14}{12}}{1-(1+\frac{.14}{12})^{-300}}\]

OpenStudy (anonymous):

i get $1567.89 per month, multiply by 12 to get yearly payment http://www.wolframalpha.com/input/?i=%28130000*.14%2F12%29%2F%281-%281%2B.14%2F12%29^%28-300%29%29